pts) Answer the following questions about two disaccharides, (a) maltose, or malt sugar, (b) lactose, or milk sugar. ОНН он н нон Maltose contains two glucose monomers Lactose contains one galactose and one glucose bond pts) Explain what is meant by a, and 4 in the term a-1,4 glycosidic bond. rcle this bond in the appropriate structure. ts) Explain what is meant by B, 1, and 4 in the term p-1,4 glycosidic bor le this bond in the appropriate structure. s) By extension, explain what would be meant by the term a.-1,6 glycosidi
Answer #1

Maltose is a sugar formed when two glucose molecules are joined together by a glycosidic bond by removing one water molecule. This is illustrated in the diagram below:

CH,OH CH2OH CH2OH CH2OH 0 0. C1 C4 + H20 OH OH OH OH OH 앤--… OH OH OH a-1,4-glycosidic OH OH OH bond OH glucose glucose maltose water

Here, two \alpha -D- glucose molecules are joined together in the same conformation. This means that the C1 of the first glucose and C4 of the second glucose are joined together by this bond. A glycosidic bond is formed between C1 and C4 when one water molecule is removed from these and a covalent bond is formed.
The \alpha here means that the OH groups from which water is removed are present below the glucose ring.
Thus, the name \alpha -1,4-glycosidic bond means that two OH groups which are present beneath the glucose molecules give away one water molecule and C1 of one glucose and C4 of the second glucose join together.

During the formation of lactose, a \beta -D galactose and \beta -D glucose join together to form a glycosidic bond known as \beta -1,4-glycosidic bond. It is also a covalent bond. Here, \beta means that the OH which forms the water molecule is present above the plane of the galactose and glucose molecules. Here also, C1 of galactose and C4 of glucose join together to form a bond and one water molecule is given out.

Lactose CH2OH CH20H HO OH OH OH OH OH

\alpha-1,6-glycosidic bond would mean that C1 of the first sugar molecule and C6 of the second sugar molecule is joined together by a covalent bond and in this process, a water molecule is given out. The OH groups which form this water molecule would be present below the plane of the sugar molecules.